A card company mixes two varieties of cards. Embossed cards cost $.65 each, and regular cards $.40 each. How many cards of each type should be included in an assortment of 25 cards that costs $14?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

A card company mixes two varieties of cards. Embossed cards cost $.65 each, and regular cards $.40 each. How many cards of each type should be included in an assortment of 25 cards that costs $14?
--------------------
Let "x" be the # of embossed cards: value is 0.65x dollars
--------------------
# of regular cards is "25-x" ; value is 0.40(25-x)=10-0.40x dollars
---------------------
EQUATION:
value + value = 14 dollars
0.65x + 10-0.40x = 14
0.25x = 4
x = 16 (# of embossed cards in the mix)
25-x = 9 (# of regular cards in the mix)
==================
Cheers,
Stan H.

Answer by checkley71(8403)

.65X+.40(25-X)=14
.65X+10-.40X=14
.25X=14-10
.25X=4
X=4/.25
X=16 $.65 CARDS WERE SOLD.
25-16=9 $.40 CARDS WERE SOLD
PROOF
.65*16+9*.40=14
10.4+3.6=14
14=14